# Democratizing Uncertainty Quantification

### v2

@@ 1 Main
These successes notwithstanding, advanced statistical inference tools are not yet employed extensively in conjunction with computationally demanding models. Such models appear, for example, in engineering design, in geophysics, or in medical imaging. We believe that this is in part due to a chicken-and-egg problem: UQ has not been demonstrated to be useful in those areas, but demonstration requires advanced knowledge of both application and UQ techniques that few research groups have. Our work addresses both of those issues by reducing the required level of knowledge via generic interfaces.
@@
Its unified interface allows linking arbitrary tools and makes it easy to swap out components on either side, thus reducing user lock-in. UM-Bridge further offers portable models through optional containerization, which offers a high degree of separation of concerns between UQ and model experts.
In order to strengthen rigorous performance comparisons in UQ, we present what we believe to be the first library of UQ benchmark problems, developed by the authors of this paper and provided as part of an open-source software repository. UM-Bridge support makes the benchmarks available to virtually any UQ software and ensures portability and reproducibility via containerization. We describe these UQ benchmarks in detail in Appendix B.
In addition, UM-Bridge offers a universal approach to scale-up UQ applications on parallel compute clusters that is fully model agnostic. In a black-box fashion, UM-Bridge even allows for HPC-scale computations with existing UQ codes that had never been intended for that purpose. In Appendix A, we describe in detail the workflow for three complex UQ applications, each involving the coupling of UQ software with HPC code.
Thus, UM-Bridge democratizes UQ and paves the way to its widespread adoption by seperating concerns of UQ, model and HPC experts, providing transparent scalability and facilitating reproducible benchmarking.
2 Results
@@ 2.1 Making tools accessible
Figure 2: Fortran code for modelling the pressure distribution and the wave elevation generated by a vessel. UM-Bridge support is achieved through a simple wrapper, enabling parallelized UQ from, e.g., Matlab.
More generally, any language supporting HTTP communication can implement UM-Bridge support. Models that are written in unsupported languages or that cannot be embedded in a higher-level code may still reap most benefits of UM-Bridge when called from a simple UM-Bridge wrapper. As an example that demonstrates how previously incompatible applications are now readily available to high-level UQ tools, in Section A.1 we seamlessly apply a Matlab UQ code to the Fortran-written numerical ship hull model [20] in Figure 2.
@@ 2.2 Separating concerns
Building UQ applications requires expertise in both UQ and the numerical model at hand. In addition to a theoretical understanding, it also calls for experience with software packages from both domains. Such an overarching expertise is often time-consuming to achieve or outright infeasible within the constraints of a project.
@@
Figure 3: FE model of a carbon fibre composite aerospace component implemented in a complex PDE solver in C++. Through UM-Bridge and containerization, a UQ expert could solve uncertainty propagation without technical knowledge of the model code.
@@
Illustrating separation of concerns, in application Section A.2 a simulation expert developed an advanced modelling application in C++ (see Figure 3), containerized it, and shared it with a UQ expert. The UQ expert was able to quantify the parametric uncertainty using a high-level UQ package written in Python, without further integration work or any deeper technical knowledge of the model implementation.
@@ 2.3 Providing transparent scalability
We illustrate these concepts with the application in Appendix A.3. It shows how an existing UQ code can transparently offload costly model evaluations to a remote cluster of 2800 processor cores, running 100 instances of a sophisticated, parallel tsunami model.
@@ 2.4 Facilitating reproducible benchmarking
A universal interface like UM-Bridge can address this problem. Thus, we have collected a large number of benchmark problems (described in more detail in Section 4.5 below), with the goal of being:
@@
  • Easy to use: All benchmark problems have an easily accessible documentation and are implemented with UM-Bridge support. They are therefore readily available to virtually any UQ package or prototype code.
@@
Appendix B gives detailed documentation of all currently available models and benchmark problems. We intend for this library to be a growing collection and encourage future contributions to it. This collection can also serve related areas, like optimization, and an example in this direction can be found in Section B.4.
3 Discussion
Package
Link
Scalability
Multiplicity
Model language agnostic
Technology
hIPPYlib-MUQ [150]
one-to-one
No
C++, Python
Single node
MUQ [149]
one-to-many
No
C++, Python
HPC (MPI)
PyMC [7], QMCPy [109]
one-to-many
No
Python
Single node
Dakota [23]
one-to-many
Yes
C++, Files
HPC (SLURM)
lagun [24]
one-to-many
Yes
Files
HPC (SLURM)
UQLab [25]
one-to-many
Yes
Matlab, files
HPC (SLURM)
UM-Bridge [26]
many-to-many
Yes
C++, Julia, Matlab, Python, R
HPC (SLURM, PBS), Kubernetes
Limitations
Outlook
4 Methods
@@
Bayesian inference in turn defines the probability of a parameter $\theta$ given observed data $y$ in terms of a prior distribution $\pi_{\text{prior}}$ and a likelihood function $L$:
@@
The first factor on the right represents prior information about the parameter distribution we seek, while the likelihood $L(y|\theta)$ is typically defined based on a measure of distance between model prediction $F(\theta)$ and data $y$. The goal is then to characterize $\pi_{\text{posterior}}$, for example through sampling or by estimating its moments.
@@ 4.1 UM-Bridge architecture
Many forward UQ methods (e.g. many Monte Carlo (MC) samplers, or stochastic collocation [35]) only require model evaluations $\{F(\theta_{i})\}$ at an (at least in practice) finite number of points $\theta_{1},\ldots,\theta_{N}$. Other methods may additionally need derivatives of $F$ at $\theta_{i}$. Similarly, inversion problems are often solved by pointwise evaluations of $\pi_{post}$, for which different methods can be employed (Metropolis-Hastings MCMC [36], Hamiltonian MCMC [37], NUTS [4], etc.). Again, this implies evaluating the forward model $F(\theta_{i})$ at points $\theta_{i}$ and, in some cases, also derivatives.
Generally, across UQ algorithms, the model is often assumed to provide some of the following pointwise mathematical operations:
  • Evaluation $F(\theta)$,
@@
  • Jacobian action $J_{F}(\theta)v$, and
@@
Figure 4: Monolithic coupling of UQ and model software in a single application (top); UM-Bridge providing a universal interface between UQ and model applications (bottom). UM-Bridge integrations (”UM”) handle network communication behind the scenes. Optionally, models may be containerized.
The key idea behind UM-Bridge is to implement this mathematical interface in an equally universal software interface: We treat UQ algorithm and numerical model as stand-alone applications, linked only through an Hypertext Transfer Protocol (HTTP) based network protocol (see Figure 4, [26]). In contrast to a monolithic approach, this microservice-inspired architecture enables linking across arbitrary languages and software packages, as well as separation of concerns. It should be noted that multilevel, multi-index or multi-fidelity methods (e.g. [38, 39, 113, 41, 42, 114, 44, 45, 46, 47]) such as the one in Section A.3 operate on an entire hierarchy of models. Less accurate but cheaper-to-compute models are then used to accelerate the UQ method, leading to considerable gains in efficiency both in forward and inverse UQ problems. While this necessitates multiple models, each individual model still fits in the framework described above.
@@ 4.2 The UM-Bridge protocol
UM-Bridge treats model inputs and outputs as lists of vectors, which is inspired by MUQ’s proven internal model interface [149]. This enables automatic differentiation with respect to certain components of the input, even when chaining multiple models.
The aforementioned integrations (see Table 1) implement the protocol behind the scenes, making UM-Bridge models available either as simple function calls or fully integrated in the respective software package’s model structure.
@@ 4.3 Basic examples
Let us illustrate how UM-Bridge works in practice with some simple examples. While intentionally basic, the same approach underlies the applications in Appendix A.
@@ 4.3.1 Clients - Interacting with models
First, we need a running model server, providing a forward operator $F$. This could be a simulation code with UM-Bridge support running directly on your system, a custom one as in Section 4.3.2, or a containerized model. For example, we could download and run the containerized tsunami model from the benchmark library (see Section B.1.3) via docker:
@@
All that is needed here is the URL at which the model server is running, here http://localhost:4242=, the name of the requested model $F$, here and a parameter $\theta$ to evaluate $F$ at. The actual evaluation is then a call to the model with a given parameter $\theta=(0,10)$, which returns $F(\theta)$. Optionally, we can pass arbitrary model-specific configuration options, for example to set the discretization level of a multifidelity model:
@@ 4.3.2 Servers - Defining models
The code sets The final line instructs UM-Bridge to make the model available to clients, acting as a server. UM-Bridge servers written in other languages follow the same principle.
@@ 4.4 Scaling up on clusters
As the client-side interface is entirely identical, a UQ software is completely oblivious to the model being executed on a cluster instead of the local machine. However, the UQ software may now send multiple concurrent evaluation requests to the cluster. Since the model evaluation is typically costly compared to the UQ algorithm itself, a UQ package spawning hundreds of threads on a laptop offloading work to a large cluster is perfectly viable.
Likewise, since every model instance only receives sequential model evaluations via UM-Bridge, no modification to the model is needed compared to running on a single machine. In addition, containerized models can be run on cloud or HPC systems without additional setup cost as long as the HPC systems supports containers.
Since the UQ software is not involved in distributing evaluations across the cluster, parallelism within many existing UQ packages works out of the box with this setup. For example, the Sparse Grids Matlab Kit (SGMK), MUQ, QMCPy, and PyMC have all been tested successfully across a number of different models. Some of these applications are discussed in Appendix A.
Note that, due to the flexibility of the interface, more complex architectures are easy to construct. For example, as in Section A.3, a multilevel UQ code may at the same time be pointed to a fast surrogate model running on the same workstation and a remote cluster for full-scale model runs, accommodating for heterogeneous resource demand across levels
@@ 4.4.1 UM-Bridge on kubernetes clusters
The only changes needed for a custom application are docker image name, number of instances (replicas) and resource requirements in the provided model.yaml configuration. Below is an example of an adapted model configuration, namely the one used in the sparse grids application in Section A.1:
@@
Section A.3 demonstrates the UM-Bridge kubernetes setup on 2,800 processor cores running 100 parallel model instances. We further performed synthetic weak scaling tests running up to $N=512$ instances of the L2-Sea model (see Section B.1.2), configured to use 1 processor core each and taking around 2.5s per evaluation. We then requested $20\cdot N$ model evaluations from the cluster. As Figure 6 shows, we did not observe a significant bottleneck, indicating that this setup could be scaled further. In addition, the model itself could be parallelized. With a model parallelized across 100 cores, we expect similar results to the 512 instance test while fully utilizing a 51,200 core cluster.
@@ 4.5 UQ Benchmark library
Section B.1 contains detailed descriptions of all models, Section B.2 describes inference benchmarks and Section B.3 describes propagation problems. Tables 4.5.1, 4, and 5 provide brief summaries.
4.5.1 Definition of models
The models in our library range from simple analytic functions to complex numerical simulations. In contrast to the propagation benchmarks, models intentionally offer configuration options, allowing users to explore model variations and to adapt them to their needs. For example, the tsunami model can be set to operate on three different refinement levels. When developing algorithms, operating on fast approximate model levels may be more suitable for quick test runs.
@@ 4.5.1 Definition of models
{xtabular}p0.24p0.64 Euler-Bernoulli beam & Deformation of an Euler-Bernoulli beam with a spatially variable stiffness parameter.
@@
4.5.2 Bayesian inference benchmarks
Inference benchmarks define a (potentially not normalized) probability density function $\pi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ whose properties are to be estimated. They are therefore additionally restricted to a one-dimensional output. For numerical stability, our definition of the interface requires implementations to provide the logarithm of the $\pi$.
Some of the benchmarks (see previous sections) also define a Bayesian inference problem based on a model and fixed data $y$,
@@ 4.5.2 Bayesian inference benchmarks
[…]
Name Short description
[…]
{xtabular}p0.24p0.64 Analytic densities Infers the PDF of various analytic functions.
@@
4.5.3 Uncertainty propagation benchmarks
@@
Propagation Benchmarks
Euler-Bernoulli beam
Modelling the effect of uncertain material parameters on the displacement of an Euler-Bernoulli beam with a prescribed load.
Genz
Multi-dimensional functions for testing quadrature and surrogate methods.
L2-Sea UQ
Forward UQ of the total resistance in calm water conditional to operational and geometrical uncertain parameters.
Cookies problem
Forward UQ of average temperature in a corner of the “oven”
The goal of uncertainty propagation is to determine the effect of uncertainties in model parameters on model outputs. The parameter $\theta$ is specified to be a random variable with a distribution documented in the markdown file. The solution of a propagation benchmark is then an approximation to the distribution of $F(\theta)$, or an estimate of some derived quantity like $\mathbb{E}[G(F(\theta))]$ where $G$ is an “output functional”.
4.5.4 Optimization benchmarks
UM-Bridge’s interface is fundamentally agnostic to what algorithms do with what the interface provides. This makes it possible to also use UM-Bridge for black-box optimization, and we have included an example of an optimization benchmark, see Table 6 and Section B.4.

### v4

@@ Abstract
keywords:
Uncertainty Quantification , Numerical Simulation , Benchmarks , Scientific Software , High-Performance Computing
journal: Computational Physics
[…]
@@ 1 Introduction
Notwithstanding those successes, advanced statistical inference tools are not yet employed extensively in conjunction with computationally demanding models. Such models appear, for example, in engineering design, in geophysics, or in medical imaging. We believe that this is in part due to a chicken-and-egg problem: UQ has not been demonstrated to be useful in those areas, but demonstration requires advanced knowledge of both application and UQ techniques that few research groups have. Our work addresses both of those issues by reducing the required level of knowledge via generic interfaces.
@@
Its unified interface allows linking arbitrary tools and makes it easy to swap out components on either side, thus reducing user lock-in. UM-Bridge further offers portable models through optional containerization, which offers a high degree of separation of concerns between UQ and model experts. In order to strengthen rigorous performance comparisons in UQ, we present what we believe to be the first library of UQ benchmark problems, developed by the authors of this paper and provided as part of an open-source software repository. UM-Bridge support makes the benchmarks available to virtually any UQ software and ensures portability and reproducibility via containerization. We describe these UQ benchmarks in detail in Appendix A.
In addition, UM-Bridge offers a universal approach to scale-up UQ applications on parallel compute clusters that is fully model agnostic. In a black-box fashion, UM-Bridge allows for HPC-scale computations with UQ codes that had never been intended for that purpose. In Section 5, we describe in detail the workflow for three complex UQ applications, where UQ software is successfully coupled with an HPC code.
Thus, by seperating concerns of UQ, model and HPC experts, UM-Bridge democratizes UQ and paves the way to its widespread adoption, providing transparent scalability and facilitating reproducible benchmarking.
2 UM-Bridge – a new paradigm for UQ software integration
@@ 2.1 Making tools accessible
More generally, any language supporting HTTP communication can implement UM-Bridge support. Models that are written in unsupported languages or that cannot be embedded in a higher-level code may still reap most benefits of UM-Bridge when called from a simple UM-Bridge wrapper. As an example that demonstrates how previously incompatible applications are now readily available to high-level UQ tools, in Section 5.1 we seamlessly apply a Matlab UQ code to a Fortran-based model of a ship hull [20], as shown in Figure 2 (left).
Figure 2: Fortran code for modelling the pressure distribution and the wave elevation generated by a vessel. UM-Bridge support is achieved through a simple wrapper, enabling parallelized UQ from, e.g., Matlab (left figure). FE model of a carbon fibre composite aerospace component implemented in a complex PDE solver in C++. Through UM-Bridge and containerization, a UQ expert could solve uncertainty propagation without technical knowledge of the model code (center & right figures).
@@ 2.2 Separating concerns
Building UQ applications requires expertise in both UQ and numerical modelling. In addition to a theoretical understanding, it also calls for experience with software packages in both domains. It is often time-consuming or outright infeasible to achieve uch an overarching expertise the constraints of a project.
@@
Illustrating separation of concerns, in application Section 5.2 a simulation expert developed an advanced modelling application in C++, containerized it, and shared it with a UQ expert, see Figure 2 (right). The UQ expert was able to quantify the parametric uncertainty using a high-level UQ package written in Python, without further integration work or any deeper technical knowledge of the model implementation.
@@ 2.3 Providing transparent scalability
We illustrate these concepts with the application in Appendix 5.3. It shows how an existing UQ code can transparently offload costly model evaluations to a remote cluster of 2800 processor cores, running 100 instances of a sophisticated, parallel tsunami model.
@@ 2.4 Facilitating reproducible benchmarking
A universal interface like UM-Bridge can address this problem. Thus, we have collected a large number of benchmark problems (described in more detail in Section 4 below), with the goal of being:
  • 1.
@@
  • 2.
    Easy to use: All benchmark problems are implemented with UM-Bridge support and have easily accessible documentation. They can readily be used to benchmark any UQ package or prototype code.
  • 3.
@@
Appendix A gives detailed documentation of all currently available models and benchmark problems. We intend for this library to be a growing collection and encourage future contributions to it. This collection can also serve related areas, like optimization, and an example in this direction can be found in Section A.4.
3 UM-Bridge architecture and usage
@@
On the other hand, Bayesian inference defines the probability of a parameter $\theta$ given observed data $y$ in terms of a prior distribution $\pi_{\text{prior}}$ and a likelihood function $L$:
@@
The first factor on the right represents prior information about the parameter distribution we seek, while the likelihood $L(y|\theta)$ provides a measure of distance between model prediction $F(\theta)$ and data $y$. The goal is then to characterize $\pi_{\text{posterior}}$, for example through sampling or by estimating its moments.
@@ 3.1 UM-Bridge architecture
Many forward UQ methods (e.g. many Monte Carlo (MC) samplers, or stochastic collocation [22]) only require model evaluations $\{F(\theta_{i})\}$ at a finite number of points $\theta_{1},\ldots,\theta_{N}$. Other methods may additionally need derivatives of $F$ at $\theta_{i}$. Similarly, inversion problems are often solved by pointwise evaluations of $\pi_{post}$, for which different methods can be employed (Metropolis-Hastings MCMC [23], Hamiltonian MCMC [24], NUTS [4], etc.). Again, these approaches require only forward model evaluations $F(\theta_{i})$ and, in some cases, derivatives at a finite number of points $\theta_{1},\ldots,\theta_{N}$.
In general, across most UQ algorithms, the model is typically assumed to provide (some of) the following pointwise mathematical operations:
  • 1.
    Model evaluation $F(\theta)$,
  • 2.
@@
  • 3.
    Jacobian action $J_{F}(\theta)v$,
  • 4.
@@
Figure 3: Monolithic coupling of UQ and model software in a single application (left); UM-Bridge providing a universal interface between UQ and model applications (right). UM-Bridge integrations (”UM”) handle network communication behind the scenes. Optionally, models may be containerized.
The key idea behind UM-Bridge is to implement the above mathematical interface in an equally universal software interface: We treat UQ algorithm and numerical model as stand-alone applications, linked only through an Hypertext Transfer Protocol (HTTP) based network protocol (see Figure 3, [25]). In contrast to a monolithic approach, this microservice-inspired architecture enables linking across arbitrary languages and software packages, as well as separating concerns. It should be noted that multilevel, multi-index or multi-fidelity methods (e.g. [26, 27, 28, 29, 30, 31, 32, 33, 34, 35]) such as the one in Section 5.3 operate on an entire hierarchy of models. Less accurate but cheaper-to-compute models are used to accelerate the UQ method, leading to considerable gains in efficiency both in forward and inverse UQ problems. While this necessitates multiple models, each individual model still fits in the framework described above.
@@ 3.2 The UM-Bridge protocol
UM-Bridge treats model inputs and outputs as lists of vectors, which is inspired the proven internal model interface in MUQ [12]. This enables automatic differentiation with respect to certain components of the input, even when chaining multiple models.
In all the aforementioned integrations of UM-Bridge (see Table 1), the protocol is implemented behind the scenes, making UM-Bridge models available either as simple function calls or fully integrated in the respective software package’s model structure.
@@ 3.3 Basic examples
Let us illustrate how UM-Bridge works in practice with some simple examples. While intentionally basic, the same approach underlies the applications in Section 5.
@@ 3.3.1 Clients - Interacting with models
First, we need a running model server, providing a forward operator $F$. This could be a simulation code with UM-Bridge support running directly on your system, a custom one as in Section 3.3.2, or a containerized model. For example, we could download and run the containerized tsunami model from the benchmark library (see Section A.1.3) via docker:
@@
All that is needed is the URL of the model server (here http://localhost:4242), the name of the requested model $F$ (here forward) and a parameter $\theta$ to evaluate $F$ at. The actual evaluation is then a call to the model with a given parameter $\theta=(0,10)$, which returns $F(\theta)$. Optionally, arbitrary model-specific configuration options can be passed, for example to set the discretization level of a multifidelity model:
@@ 3.3.2 Servers - Defining models
The code sets supports_evaluate to indicate to UM-Bridge that model evaluations are supported. Likewise, gradients, Jacobian- or Hessian-action could be implemented and marked as supported.
The final line instructs UM-Bridge to make the model available to clients, acting as a server. UM-Bridge servers written in other languages follow the same principle.
@@ 3.4 Scaling up on clusters
As the client-side interface is entirely identical, a UQ software is completely oblivious to the model being executed on a cluster instead of the local machine. However, the UQ software may now send multiple concurrent evaluation requests to the cluster. Since the model evaluation is typically costly compared to the UQ algorithm itself, a UQ package spawning hundreds of threads on a laptop offloading work to a large cluster is perfectly viable. Likewise, since every model instance only receives sequential model evaluations via UM-Bridge, no modification to the model is needed compared to running on a single machine. In addition, containerized models can be run on cloud or HPC systems without additional setup cost as long as the HPC systems supports containers.
Since the UQ software is not involved in distributing evaluations across the cluster, parallelism within many existing UQ packages works out of the box with this setup. For example, the Sparse Grids Matlab Kit (SGMK), MUQ, QMCPy, and PyMC have all been tested successfully across a number of different models. Some of these applications are discussed in Section 5.
Note that, due to the flexibility of the interface, more complex architectures are easy to construct. For example, as in Section 5.3, a multilevel UQ code may at the same time be pointed to a fast surrogate model running on the same workstation and a remote cluster for full-scale model runs, accommodating for heterogeneous resource demand across levels
@@ 3.4.1 UM-Bridge on kubernetes clusters
The only changes needed for a custom application are docker image name, number of instances (replicas) and resource requirements in the provided model.yaml configuration. Below is an example of an adapted model configuration, namely the one used in the sparse grids application in Section 5.1:
@@
Section 5.3 demonstrates the UM-Bridge kubernetes setup on 2,800 processor cores running 100 parallel model instances. We further performed synthetic weak scaling tests running up to $N=512$ instances of the L2-Sea model (see Section A.1.2), configured to use 1 processor core each and taking around 2.5s per evaluation. We then requested $20\cdot N$ model evaluations from the cluster. As Figure 5 shows, we did not observe a significant bottleneck, indicating that this setup could be scaled further. In addition, the model itself could be parallelized. With a model parallelized across 100 cores, we expect similar results to the 512 instance test while fully utilizing a 51,200 core cluster.
@@ 4 UQ Benchmark library
Section A.1 contains detailed descriptions of all models, Section A.2 describes inference benchmarks and Section A.3 describes propagation problems. Tables 2, 3, and 4 provide brief summaries.
4.1 Models
@@
Euler-Bernoulli beam Deformation of an Euler-Bernoulli beam with a spatially variable stiffness parameter.
@@
The models in our library range from simple analytic functions to complex numerical simulations (see Table 2). In contrast to the propagation benchmarks, models intentionally offer configuration options, allowing users to explore model variations and to adapt them to their needs. For example, the tsunami model can be set to operate on three different refinement levels. When developing algorithms, operating on fast approximate model levels may be more suitable for quick test runs.
4.2 Bayesian inference benchmarks
Inference benchmarks define a (potentially not normalized) probability density function $\pi:\mathbb{R}^{n}\rightarrow\mathbb{R}$ whose properties are to be estimated. They are therefore restricted to a one-dimensional output. For numerical stability, our definition of the interface requires implementations to provide the logarithm of $\pi$. See Table 3 for a list of inference problems that are currently part of the library.
Some of the benchmarks define a Bayesian inference problem based on a model and fixed data $y$,
@@ 4.2 Bayesian inference benchmarks
Analytic densities Infers the PDF of various analytic functions.
@@
4.3 Uncertainty propagation benchmarks
@@
Euler-Bernoulli beam Modelling the effect of uncertain material parameters on the displacement of an Euler-Bernoulli beam with a prescribed load.
Genz Multi-dimensional functions for testing quadrature and surrogate methods.
L2-Sea UQ Forward UQ of the total resistance in calm water conditional to operational and geometrical uncertain parameters.
Cookies problem Forward UQ of average temperature in a corner of the “oven”
The goal of uncertainty propagation is to determine the effect of uncertainties in model parameters on model outputs. The parameter $\theta$ is specified to be a random variable with a distribution documented in the markdown file. The solution of a propagation benchmark is then an approximation to the distribution of $F(\theta)$, or an estimate of some derived quantity like $\mathbb{E}[G(F(\theta))]$ where $G$ is an “output functional”. Table 4 shows the currently available propagation benchmarks.
4.4 Optimization benchmarks
UM-Bridge’s interface is fundamentally agnostic to what algorithms do with what the interface provides. This makes it possible to also use UM-Bridge for other fields like black-box optimization. We have included an initial example of an optimization benchmark in Table 5 and Section A.4.